An engineer planning for his child’s college education purchased a zero-coupon corporate bond (i.e. a bond that has no dividend payment) fro $7500. The bond has a face value of $50,000 and is due in 18 years. If the bond is held to maturity, determine the ROR, i* for the investment. (Enter answer in 2 decimal places. If your answer is 8.25%, enter 8.25).
Answer #1

Formula is F = P *(1+i)^t

F = 50000, P = 7500, t = 18 yrs

50000 = 7500 * (1+i)^18

(1+i)^18 = 50000 / 7500 = 6.666667

1+i = 6.666667 ^ (1/18) = 1.11115

i = 1.11115 - 1 = 0.11115

i = 0.11115 = 11.11 %